# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v7

@@ 10.1 Introduction
So far, this battle has been a battle between a catalyst, the chosen minimum degree node, and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
@@
Let $G$ be an oriented graph represented by a GLOVER, with all nodes satisfying the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. If there are back arcs in $G$, then a node’s degree will double.
@@ 10.2 Theorem Statement
Let $G$ be an oriented graph represented by a GLOVER with all nodes having the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node where we set up our partition. Then Seymour’s conjecture holds true.
@@
We have set up the containers using a GLOVER data structure. There are no back arcs; else we will have a degree-doubling node. Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
@@
In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, one must understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So, for small graphs, we are just not going out far enough to catch up. For larger graphs, though, we do start to see this.
@@
Occurrence of Last Dense Container Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$. Then there will be a container that reaches maximal density. Further, if that container is $C_{i}$, it will be $i$-out-regular.
@@
Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none. At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively getting to a denser container. As described in corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever. There must be a last, most dense container. ∎
@@
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
@@
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group. Dense regular graphs have many applications in themselves, from areas like expander graphs to network topology design to coding theory.
@@
First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles. Secondly, we also note that the lemma 6.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container. Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another. The combination of these three lemmas lets us know that every node inside a container has their interior degree doubled. This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
@@
We opened by showing several applications of the SSNC, including social media, social network modeling and network analysis. This highlights the necessity for an algorithm that can work efficiently in practice. This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph is important. This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively. That makes it potentially applicable to real-world networks with millions or billions of nodes and edges. This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
@@ 11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms, and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
@@
The SSNC started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
@@
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees.
@@
Finally, this research offered an alternative method for the SSNC. We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists. To do this, we developed an effective algorithm and proved its complexity. Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions. This may encourages researchers to explore constructive approaches with a focus on practical applications. Hopefully, this work will inspire further research on constructive approaches to other graph theory problems. Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
@@ 14 Appendix 2: Applicability of the Graph Level Order
What makes GLOVER special are things like total order on containers. Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting. This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance. This is hard to measure on a single scale.
@@ 15 Appendix 3: Visualization and Accessibility of Oriented Graphs
Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com. This is an evolving project. Additional examples and applications will continue to be added. Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement. Our goal here was to help users connect with the SSNC. Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science. Thank you for your engagement.

### v8

@@ 11.1 Introduction
So far, this battle has been a battle between a catalyst, $v_{0}$, and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
@@
Let $G$ be an oriented graph, with all nodes satisfying the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node in $G$. If there are back arcs in $G$, then a node’s degree will double.
@@ 11.2 Theorem Statement
Let $G$ be an oriented graph, with all nodes having the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node where we set up our partition. Then Seymour’s conjecture holds true.
@@
We have set up the containers using a GLOVER data structure. There are no back arcs; else we will have a degree-doubling node. Next is lemma 8.1 (Interior Degrees Doubling), which states that in every container we will have a node whose interior degree doubles.
@@
In small graphs, such as the one shown in Figure 4.3 (Minimum Out-Degree 3 with Neighbors 1), the collision phenomenon was not observed. This is because the container sizes decrease linearly (at a rate of 1), while the interior degrees increase linearly (also at a rate of 1). However, the number of arcs within the oriented graph grows quadratically, at a rate of $²O(n\texttwosuperior)$. Consequently, while the number of nodes increases linearly, the number of edges increases quadratically. In small graphs, the growth hasn’t progressed far enough for the quadratic edge growth to outpace the linear node growth and cause collisions. In larger graphs, however, this difference in growth rates becomes significant, and collisions begin to occur.
@@
Occurrence of Last Dense Container Let $G$ be an oriented graph, with minimum degree $k$. Then there will be a container that reaches maximal density. Further, if that container is $C_{i}$, it will be $i$-out-regular.
@@
Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree node. At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively getting to a denser container. As described in corollary 11.1 (Degree Doubling by Container Density), this cannot continue forever. There must be a last, most dense container. ∎
@@
Multiple Degree Doubling Nodes Let $G$ be an oriented graph, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
@@
Then third, we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group. Dense regular graphs have many applications in themselves, from areas like expander graphs to network topology design to coding theory.
@@
First, remember that we proved in lemma 8.1 (Interior Degrees Doubles) that every container has a node whose interior degree doubles. Secondly, we also note that the lemma 7.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container. Third, lemma 9.2 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another. The combination of these three lemmas lets us know that every node inside a container has their interior degree doubled. This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
@@
We opened by showing several applications of the SSNC, including social media, social network modeling, and network analysis. This highlights the necessity for an algorithm that can work efficiently in practice. This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph, is important. This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively. That makes it potentially applicable to real-world networks with millions or billions of nodes and edges. This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
@@ 12 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms, and data structures. This research helped us to use many mathematical tools, like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
@@
The SSNC started as a question in 1990. Over the next 34 years, it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
@@
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees. Still, the ability for the data structure to partition the nodes of an oriented graph into containers, which then allowed for powerful tools like divide-and-conquer and mathematical induction, can have further applications. This would then allow us to implement resources from the field of graph theory into subject matter dealing with machine learning, data science, and network flow.
@@
Finally, this research offered an alternative method for the SSNC. We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists. To do this, we developed an effective algorithm and proved its complexity. Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions. This may encourage researchers to explore constructive approaches with a focus on practical applications. Hopefully, this work will inspire further research on constructive approaches to other graph theory problems. Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
@@ 15 Appendix 2: Applicability of the Graph Level Order
What makes GLOVER special are things like a total order on containers. Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting. This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance. This is hard to measure on a single scale.
@@ 16 Appendix 3: Visualization and Accessibility of Oriented Graphs
Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com.
This is an evolving project. Additional examples and applications will continue to be added.
Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement.
Our goal here was to help users connect with the SSNC. Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science. Thank you for your engagement.